OpenStudy (anonymous):

In a rectangle height is one third of the base length. Area of ​​the rectangle is 108 cm2. Determine the rectangle base and height. Justify your answers with words and calculations. b= h=

OpenStudy (nirmalnema):

let the base be x...so height is x/3 area = h*b x*x/3=108 x^2/3=108 x^2=108*3 x^2=324 x= rt.ov.324 x= 18 so base = 18 cm and height = 6cm

OpenStudy (nirmalnema):

u get it?? @sarapetrini

OpenStudy (nirmalnema):

??/

OpenStudy (anonymous):

I have it like I have to write the step by step on: B: a big empty space and then separated from that H: a big empty space :S

OpenStudy (nirmalnema):

i have given it step by step

OpenStudy (anonymous):

but there is like i have to write the step by step of H and step by step of B, like 2 step by step of each of them.. aaah i write all of it together:P Thank u again for ur Huge help nirmalnema,<3<3

OpenStudy (nirmalnema):

yaa u can write it toghther.. and again welcome... <3

OpenStudy (nirmalnema):

hey i got to go now so u can take further help from .. @aayushi.somani help her with the further problems..!!

OpenStudy (anonymous):

thank u so much nirmalnema!!!!!<3

OpenStudy (nirmalnema):

no problem... <3..!!